OpenStudy (anonymous):

Solve the following system of equations. Please show your work x - y = 13 2x + y = 2

OpenStudy (anonymous):

x=13+y 2(13+y)+y=2 26+2y=2 2y=-24 y=-12 x=13+(-12) x=1

OpenStudy (anonymous):

are u used elimination for this equation ???

OpenStudy (anonymous):

or subtitution !!

OpenStudy (anonymous):

I used substitution :)